I add 3 to the number, divide the result by 5, I get the same answer as I take away 5 from it
Found 2 solutions by brysca, Cromlix:
Answer by brysca(112)

Solution posted below
.
hope this helps!

Answer by Cromlix(4381)

Hi there,
Make 'x' the number
add 3 to the number, divide the result by 5
(x + 3)/5
= the same answer as I take away 5 from it
(x + 3)/5 = x - 5
Common denominator 5
x + 3 = 5(x - 5)
x + 3 = 5x - 25
Collect like terms
x - 5x = -25 - 3
-4x = -28
x = 7
The number is 7
Hope this helps:-)